# https://sarahweiii.github.io/meshlrm/ MeshLRM: Large Reconstruction Model for High-Quality Meshes

### v1

@@ 3.2 Stage 1: Efficient Training for Volume Rendering
We train our model with ray marching-based radiance field rendering (as in NeRF [41]) to bootstrap our model, providing good initialized weights for mesh reconstruction. Instead of training directly using high-res () input images like prior LRM works [23, 30], we develop an efficient training scheme, inspired by ViT [16]. Specifically, we first pretrain our model with images until convergence and then finetune it for fewer iterations with images. This training scheme cuts the computing cost significantly, hence obtaining much better quality than training at 512-res from scratch given the same amount of compute (see Tab. 1).
@@
512-res finetuning. For high-res fine-tuning, we use 512-resolution images for the model input and output. We use a batch size of 2 per GPU and densely sample 512 points per ray, which is the number in low-res pretraining. Here, we compensate for the increased computational costs (from longer token sequences and denser ray samples) by reducing the batch size 4 times, achieving a similar training speed (per iteration) to the low-res pretraining. It’s also worth noting that our resolution-varying pretraining and finetuning scheme benefits from the use of Plücker coordinates for camera conditioning. It avoids the positional embedding interpolation [16, 46] and has inherent spatial smoothness.
Loss. We use an L2 regression loss and a perceptual loss (proposed in [9]) to supervise the renderings from both phases. Since rendering full-res images is not affordable for volume rendering, we randomly sample a patch from each target 256- or 512-res image for supervision with both losses. We also randomly sample 4096 pixels per target image for additional L2 supervision, allowing the model to capture global information beyond a single patch. The loss for volume rendering training is expressed by:
@@ 3.3 Stage 2: Stable Training for Surface Rendering
Mesh extraction and rendering. We compute a density grid by decoding the triplane features and adopt a recent differentiable marching cubes (DiffMC) technique [70] to extract mesh surface from the grid. The DiffMC module is based on a highly optimized CUDA implementation, much faster than existing alternatives [52, 51], enabling fast training and inference for mesh reconstruction.
To compute the rendering loss we need to render our generated mesh. We do so using a differentiable rasterizer Nvdiffrast [29] and utilize the triplane features to (neurally) render novel images from our extracted meshes. This full rendering process is akin to deferred shading where we first obtain per-pixel XYZ locations via differentiable rasterization before querying the corresponding triplane features and regressing per-pixel colors using the color MLP. We supervise novel-view renderings with ground-truth images, optimizing the model for high-quality end-to-end mesh reconstruction.
However, using the rendering loss alone leads to high instability during training and severe floaters in the meshes (see Fig. 3). This is mainly caused by the sparsity of density gradients during mesh rendering: unlike volume rendering which samples points throughout the entire view frustum, mesh (rasterization) rendering is restricted to surface points, lacking gradients in the empty scene space beyond the surface.
@@
where represents the ground truth surface point along the pixel ray, is randomly sampled along the ray between and camera origin, and is the volume density at ; when no surface exists for a pixel, we sample inside the object bounding box along the ray and use the far ray-box intersection as .
In essence, this loss is designed to enforce the empty space in each view frustum to contain near-zero density. Here we minimize the opacity value , computed using the ray distance from the sampled point to the surface. This density-to-opacity conversion functions as an effective mechanism for weighting the density supervision along the ray with lower loss values for points sampled closer to the surface. An alternative approach would be to directly regularize the density using an L1 loss, but we found this to lead to holes in surfaces as all points contribute the same, regardless of their distance to the surface.
Combined losses. To measure the visual difference between our renderings and ground-truth (GT) images, we use an L2 loss and a perceptual loss proposed in [9] similar to stage 1. To compute the ray opacity loss , we obtain surface points using the GT depth maps. In addition, to further improve our geometry accuracy and smoothness, we apply an L2 normal loss to supervise the face normals of our extracted mesh with GT normal maps in foreground regions. Our final loss for mesh reconstruction is
where we use , and in our experiments. Since mesh rasterization is significantly cheaper than volume ray marching, we render the images at full resolution (e.g., in our experiment) for supervision, instead of the random patches+rays used in Stage 1 volume training.
@@ 4.1 Dataset and Evaluation Protocols
Our model is trained on the Objaverse dataset [14] (730K objects) for the 1st-stage volume rendering training and is subsequently fine-tuned on the Objaverse-LVIS subset (46K objects) for the 2nd-stage surface rendering finetuning. Empirically, the Objaverse-LVIS subset has higher quality and previous work [48, 69, 13] shows that fine-tuning favors quality more than quantity. We thus follow this common practice. We evaluate the reconstruction quality of MeshLRM alongside other existing methods on the GSO [17], NeRF-Synthetic [41], and OpenIllumination [32] datasets, employing PSNR, SSIM, and LPIPS as metrics for rendering quality and bi-directional Chamfer distance (CD) as the metric for mesh geometry quality. For the CD metric, since we cannot expect the model to precisely reconstruct unseen parts, we cast rays from all test views and sample 100K points at the ray-surface intersections for each object. The Chamfer distance is computed using the points sampled from both the ground-truth and reconstructed mesh within a bounding box, measured in unit. All results are generated using four sparse-view input images.
@@ 4.2 Analysis and Ablation Study
Volume rendering (Stage 1) training strategies. To verify the effectiveness of our training strategy that uses 256-res pretraining and 512-res fine-tuning (details in Sec. 3.2), we compare with a model having the same architecture but trained with high-resolution only (i.e., 512-res from scratch). Tab. 1 shows the quantitative results on the GSO dataset with detailed training settings and timings used for the two training strategies. As seen in the table, with the same total compute budget of 64 hours and 128 GPUs, our low-to-high-res training strategy achieves significantly better performance, with a 2.6dB increase in PSNR. The key to enabling this is our fast low-res pretraining, which takes only 2.6 seconds per iteration, allowing for many more iterations to be trained within a shorter period and enabling much faster convergence. Benefiting from effective pretraining, our high-res finetuning can be trained with a smaller batch size that reduces the iteration time. In general, our 1st-stage training strategy significantly accelerates the LRM training, leading to high-quality NeRF reconstruction. This naturally improves the mesh reconstruction quality in the second stage.
@@
Effectiveness of surface fine-tuning (Stage 2). We justify the effectiveness and significance of our 2nd-stage surface fine-tuning by comparing our final meshes with those directly extracted from 1st-stage model using marching cubes on GSO dataset. The quantitative results are shown in Tab. 2, and the qualitative results are presented in the Appendix. Note that, while our 1st-stage NeRF reconstruction, i.e. MeshLRM (NeRF), can achieve high volume rendering quality, directly extracting meshes from it with Marching Cubes (MC), i.e. ‘MeshLRM (NeRF)+MC’, leads to a significant drop of rendering quality on all metrics. On the other hand, our final MeshLRM model, fine-tuned with DiffMC-based mesh rendering, achieves significantly better mesh rendering quality and geometry quality than the MeshLRM (NeRF)+MC baseline, notably increasing PSNR by 2.5dB and lowering CD by 0.58. This mesh rendering quality is even comparable — with a slight decrease in PSNR but improvements in SSIM and LPIPS — to our 1st-stage volume rendering results of MeshLRM (NeRF).
@@
Surface fine-tuning losses. We demonstrate the significance of the different losses applied during the mesh fine-tuning stage (details in Sec. 3.3) with an ablation study in Tab. 3 and Fig. 3. We observe a significant performance drop in the model without our proposed ray opacity loss, leading to severe floater artifacts as shown in Fig. 3. This is because, without supervision on empty space, the model may produce floaters to overfit to the training views rather than learning the correct geometry, as it does not receive a penalty for inaccuracies in empty space. The normal loss particularly helps to generate better geometry, i.e. a lower Chamfer distance. While the normal loss does not lead to quantitative improvements in rendering (metrics are about the same) on the GSO dataset, we still observe qualitative improvements with this loss, especially when using generated images, as shown in Fig. 3 (right). Empirically, the normal loss leads to better robustness in handling inconsistent input views, a common challenge in text-to-3D or image-to-3D scenarios (e.g. the cactus). Moreover, the geometric improvements from the normal loss are also crucial for real applications to use our meshes in 3D engines (like the one in Fig. 1), where accurate shapes are critical for high-fidelity physically-based rendering.
@@
Tiny MLPs. To justify our choices of using tiny MLPs for triplane coding, instead of a large one in previous LRMs, we compare with a version that replaces the two tiny MLPs with a 10-layer (i.e., 9 hidden layers) 64-width shared MLP decoder (as used in [23, 30]). The quantitative results on the GSO dataset are in Tab. 4. We find that the tiny MLP can achieve similar performance to the large MLP while offering a notable training speed advantage (2.7s / step vs 3.6s / step). Moreover, we observe that the heavy shared MLP hardly converge during the Stage 2 DiffMC training while the tiny separate MLPs could.
@@ 4.3 Comparison Against Baselines
In3D-LRM [30]+MC | Mesh | >60 | 1s | 500M | 1287 | 23.70 | 0.875 | 0.111 | 3.40 |
MeshLRM | >60 | 0.8s | 300M | 1283 | 27.93 | 0.925 | 0.081 | 2.68 |

### v2

@@ 3.2 Stage 1: Efficient Training for Volume Rendering
We train our model with ray marching-based radiance field rendering Mildenhall et al. (2020) to bootstrap our model, providing good initialized weights for mesh reconstruction. Instead of training directly using high-res () input images like prior LRM works (Hong et al., 2024; Li et al., 2024), we develop an efficient training scheme, inspired by ViT (Dosovitskiy et al., 2020). Specifically, we first pretrain our model with input images until convergence and then finetune it for fewer iterations with input images. This training schedule leads to significantly better quality than training at 512-res from scratch given the same amount of compute (see Tab. 1).
@@
512-res finetuning. For high-res fine-tuning, we use 512-resolution images for the model input and output. We use a batch size of 2 per GPU and densely sample 512 points per ray. Here, we compensate for the increased computational costs (from longer token sequences and denser ray samples) by reducing the batch size 4 times, achieving a similar per iteration time to the low-res pretraining. It’s also worth noting that our resolution-varying pretraining and finetuning scheme benefits from the use of Plücker coordinates for camera conditioning. It avoids the positional embedding interpolation (Dosovitskiy et al., 2020; Radford et al., 2021) and has inherent spatial smoothness.
Loss. We use an L2 regression loss and a perceptual loss (proposed in Chen & Koltun (2017)) to supervise the renderings from both phases. Since rendering full-res images is not affordable for volume rendering, we randomly sample a patch from each target 256- or 512-res image for supervision with both losses. We also randomly sample 4096 pixels per target image for additional L2 supervision, allowing the model to capture global information beyond a single patch. The loss for volume rendering training is:
@@ 3.3 Stage 2: Stable Training for Surface Rendering
Mesh extraction and rendering. We compute a density grid by decoding the triplane features and adopt a recent differentiable marching cubes (DiffMC) technique (Wei et al., 2023) to extract mesh surface from the grid. The DiffMC module is based on a highly optimized CUDA implementation, much faster than existing alternatives (Shen et al., 2023; 2021), enabling fast training and inference for mesh reconstruction.
To compute the rendering loss, we render the generated mesh using the differentiable rasterizer Nvdiffrast (Laine et al., 2020) and utilize the triplane features to (neurally) render novel images from our extracted meshes. This full rendering process is akin to deferred shading where we first obtain per-pixel XYZ locations via differentiable rasterization before querying the corresponding triplane features and regressing per-pixel colors using the color MLP. We supervise novel-view renderings with ground-truth images, optimizing the model for high-quality end-to-end mesh reconstruction.
However, using the rendering loss alone leads to high instability during training and severe floaters in the meshes (see Fig. 3). This is caused by the sparsity of density gradients in mesh rendering: unlike volume rendering which samples points throughout the entire view frustum, mesh rendering is restricted to surface points, lacking gradients in the empty scene space beyond the surface.
@@
where represents the ground truth surface point along the pixel ray, is randomly sampled along the ray between and camera origin, and is the volume density at ; when no surface exists for a pixel, we sample inside the object bounding box and use the far ray-box intersection as .
In essence, this loss is designed to encourage the empty space in each view frustum to contain near-zero density. Here we minimize the opacity value , computed using the ray distance from the sampled point to the surface. This density-to-opacity conversion functions as an effective mechanism for weighting the density supervision along the ray with lower loss values for points sampled closer to the surface. An alternative approach would be to directly regularize the density using an L1 loss, but we found this to lead to holes in surfaces as all points contribute the same, regardless of their distance to the surface. A more detailed comparison between our loss and the standard density and depth losses can be found in Appendix A.
Combined losses. To measure the visual difference between the results renderings after our DiffMC step and ground-truth (GT) images, we use an L2 loss and a perceptual loss similar to stage 1. We still supervise the opacity using , a ray opacity loss around the surface points using the GT depth maps. In addition, to further improve our geometry accuracy and smoothness, we apply an L2 normal loss to supervise the face normals of our extracted mesh with GT normal maps in foreground regions. Our final loss for mesh reconstruction is
where we use , and in our experiments. Since mesh rasterization is significantly cheaper than volume ray marching, we render the complete images ( in our experiment) for supervision, instead of the random patches+rays used in Stage 1 volume training.
@@ 4.1 Dataset and Evaluation Protocols
Our model is trained on the Objaverse dataset (Deitke et al., 2023) (730K objects) for the 1st-stage volume rendering training and is subsequently fine-tuned on the Objaverse-LVIS subset (46K objects) for the 2nd-stage surface rendering finetuning. Empirically, the Objaverse-LVIS subset has higher quality and previous work (Rombach et al., 2022; Wei et al., 2021; Dai et al., 2023) shows that fine-tuning favors quality more than quantity. We evaluate the reconstruction quality of MeshLRM alongside other existing methods on the GSO (Downs et al., 2022), NeRF-Synthetic (Mildenhall et al., 2020), and OpenIllumination (Liu et al., 2024a) datasets, employing PSNR, SSIM, and LPIPS as metrics for rendering quality and bi-directional Chamfer distance (CD) as the metric for mesh geometry quality (more details in Appendix H).
@@ 4.2 Analysis and Ablation Study
Volume rendering (Stage 1) training strategies. To verify the effectiveness of our training strategy that uses 256-res pretraining and 512-res fine-tuning (details in Sec. 3.2), we compare with a model having the same architecture but trained with high-resolution only (i.e., 512-res from scratch). Tab. 1 shows the quantitative results on the GSO dataset with detailed training settings and timings of the two training strategies. As seen in the table, with the same total compute budget of 64 hours and 128 GPUs, our low-to-high-res training strategy achieves significantly better performance, with a 2.6dB increase in PSNR. The key to enabling this is our fast low-res pretraining, which takes only 2.6 seconds per iteration with a larger batch size, allowing to train on more data within a shorter period of time and enabling much faster convergence. Benefiting from effective pretraining, our high-res finetuning can be trained with a smaller batch size that reduces the iteration time. In general, our 1st-stage training strategy significantly accelerates the LRM training, leading to high-quality NeRF reconstruction. This naturally improves the mesh reconstruction quality in the second stage.
@@
Effectiveness of surface fine-tuning (Stage 2). We justify the effectiveness and importance of our 2nd-stage surface fine-tuning by comparing our final meshes with those directly extracted from the 1st-stage model using marching cubes on the GSO dataset. The quantitative results are shown in Tab. 2, and the qualitative results are presented in the Appendix F. Note that, while our 1st-stage NeRF reconstruction, i.e. MeshLRM (NeRF), can achieve high volume rendering quality, directly extracting meshes from it with Marching Cubes (MC), i.e. ‘MeshLRM (NeRF)+MC’, leads to a significant drop of rendering quality on all metrics. On the other hand, our final MeshLRM model, fine-tuned with DiffMC-based mesh rendering, achieves significantly better mesh rendering quality and geometry quality than the MeshLRM (NeRF)+MC baseline, notably increasing PSNR by 2.5dB and lowering CD by 0.58. This mesh rendering quality is even comparable — with a slight decrease in PSNR but improvements in SSIM and LPIPS — to our 1st-stage volume rendering results of MeshLRM (NeRF).
Surface fine-tuning losses. We demonstrate the importance of the different losses applied during the mesh fine-tuning stage (details in Sec. 3.3) with an ablation study in Tab. 3 and Fig. 3. We observe a significant performance drop in the model without our proposed ray opacity loss, leading to severe floater artifacts as shown in Fig. 3. This is because, without supervision on empty space, the model may produce floaters to overfit to the input views rather than generating the correct geometry, as it does not receive a penalty for inaccuracies in empty space. Our normal loss helps to generate better geometry, i.e. a lower Chamfer distance. While it does not lead to significant quantitative improvements in rendering on the GSO dataset, we still observe qualitative improvements with this loss, especially when using generated images, as shown in Fig. 3 (right). Empirically, the normal loss leads to better robustness in handling inconsistent input views, a common challenge in text-to-3D or image-to-3D scenarios. Moreover, the geometric improvements from the normal loss are also crucial for real applications to use our meshes in 3D engines (like the one in Fig. 1), where accurate shapes are critical for high-fidelity physically-based rendering.
@@
Tiny MLPs. To justify our choice of using tiny MLPs for triplane decoding rather than the larger ones used in previous LRMs, we compare with a version that replaces the two tiny MLPs with a 10-layer (i.e., 9 hidden layers) 64-width shared MLP decoder (as used in (Hong et al., 2024; Li et al., 2024)). We show the quantitative results on the GSO dataset in Tab. 4. We find that the tiny MLP achieves similar performance to the large MLP while offering a notable training speed advantage (2.7s / step vs 3.6s / step, a 25% speedup). Moreover, we observe that the heavy shared MLP does not converge well during Stage 2 DiffMC training, unlike our proposed tiny separate MLPs.
Removal of DINO. We evaluated the benefits of DINO features and trained a MeshLRM model using the DINO encoder for the same number of steps as our proposed models. Despite having 85M (28) more pre-trained parameters than the 300M in MeshLRM, the use of DINO features results in similar performance, with only a 0.1 dB difference in PSNR and a 0.001 difference in SSIM compared to our proposed model. We also found that using only Plücker Rays for positional encoding improves our model’s zero-shot ability when processing higher resolutions (see the comparison in Appendix Fig. 8). We believe that this facilitates our high-res fine-tuning and leads to more efficient training.
@@ 4.3 Comparison Against Baselines
Comparisons with feed-forward methods.
We evaluate MeshLRM and other methods for 3D object reconstruction from four sparse views. As reported by Instant3D, previous cost-volume-based feed-forward methods (like SparseNeus(Long et al., 2022)) cannot handle this challenging sparse views case. We therefore focus on comparing with the LRM model in Instant3D (labeled as In3D-LRM) on the GSO dataset, showing quantitative results with detailed training settings and timings in Tab. 5. In the table, we show our results from both stages and compare them with both In3D-LRM and In3D-LRM + MC; for In3D-LRM + MC, Marching Cubes is directly applied to extract meshes from their NeRF reconstruction.
As shown in Tab. 5, our model, across both stages, achieves higher quality compared to In3D-LRM. As expected, In3D-LRM + MC leads to much worse rendering quality than In3D-LRM, due to the quality drop caused by the meshing process. Thanks to our effective 2nd-Stage DiffMC-based training, MeshLRM can reconstruct significantly better meshes in terms of both rendering quality (with 4.2dB PSNR and 0.05 SSIM improvements) and geometry quality (lowering the CD by 0.72). Our mesh rendering is even better than previous work volume rendering (In3D-LRM), notably increasing in PSNR by 1.4dB and SSIM by 0.03, while achieving substantially faster rendering speed. We show a qualitative comparison in Fig. 4 in which can see that our mesh renderings reproduce more texture and geometric details than the baselines. From Tab. 5, we can also see that our high reconstruction quality is achieved with smaller model size and substantially less compute, less than half of the total GPUDay compute cost required for In3D-LRM. This is enabled by our simplified LRM architecture designs and efficient low-res-to-high-res 1st-stage training strategy, leading to significantly faster training speed and convergence. Besides, our model also leads to a faster inference speed, enabling high-quality mesh reconstruction within 1 second. Overall, our MeshLRM leads to state-of-the-art sparse-view mesh reconstruction quality with high parameter utilization efficiency, and faster training, inference, and rendering speeds. We note that our simplified MeshLRM (NeRF) also shows strong performance as a state-of-the-art LRM model for sparse-view NeRF reconstruction.
@@
| In3D-LRM (Li et al., 2024)+MC | Mesh | 60 | 1s | 500M | 1287 | 23.70 | 0.875 | 0.111 | 3.40 |
| MeshLRM | 60 | 0.8s | 300M | 1283 | 27.93 | 0.925 | 0.081 | 2.68 | |
Comparisons with per-scene optimization methods.
We also compare our MeshLRM with recent per-scene optimization methods that are designed for sparse-view NeRF reconstruction, including FreeNeRF (Yang et al., 2023) and ZeroRF (Shi et al., 2023b). Our feed-forward method achieves better or comparable performance to methods that require minutes or hours to reconstruct a single scene, while taking only 0.8 seconds. Additional quantitative results and details are provided in Appendix B.
5 Applications
Our approach achieves efficient and high-quality mesh reconstruction within one second, thus facilitating various 3D generation applications when combined with multi-view image techniques.
Image-to-3D Generation
| Method | GSO | Omni3D | ||||||
| PSNR | SSIM | LPIPS | CD | PSNR | SSIM | LPIPS | CD | |
| TripoSR (Tochilkin et al., 2024) | 19.85 | 0.753 | 0.265 | 27.48 | 17.68 | 0.745 | 0.277 | 28.69 |
| LGM (Tang et al., 2024) | 18.52 | 0.713 | 0.349 | 44.41 | 14.75 | 0.646 | 0.455 | 63.38 |
| InstantMesh (Xu et al., 2024) | 20.89 | 0.775 | 0.218 | 20.80 | 17.61 | 0.742 | 0.275 | 30.20 |
| MeshFormer (Liu et al., 2024c) | 21.47 | 0.793 | 0.201 | 16.30 | 18.14 | 0.755 | 0.266 | 25.91 |
| Ours | 21.31 | 0.794 | 0.192 | 15.74 | 18.10 | 0.759 | 0.257 | 25.51 |
We follow the setup in the concurrent work MeshFormer (Liu et al., 2024c) to evaluate all methods (details in Appendix I), including TripoSR (Tochilkin et al., 2024), LGM (Tang et al., 2024), InstantMesh (Xu et al., 2024) and MeshFormer (Liu et al., 2024c). Specifically, for InstantMesh, MeshFormer, and our method, we first used Zero123++ (Shi et al., 2023a) to convert the input single-view image into multi-view images before performing 3D reconstruction. The other baselines follow their original configurations, taking a single-view image directly as input. The quantitative comparison results on GSO (Downs et al., 2022) and OmniObject3D (Wu et al., 2023) are shown in Tab. 6.
Our method demonstrates a significant advantage over TripoSR, LGM, and InstantMesh across all four metrics on both datasets. Additionally, it is comparable with MeshFormer, with slightly lower PSNR and better SSIM, LPIPS and CD. Note that MeshFormer is specifically designed for the single-image-to-3D task, it requires multi-view consistent normals as input during its reconstruction step. This makes it unsuitable for handling sparse-view reconstruction scenarios using captured images, such as setups with OpenIllumination, where only RGB images with camera poses are available. In contrast, our method can well handle this as shown in Tab. 7.
Additional visualization results are provided in Appendix Fig. 6, demonstrating that our generated meshes exhibit significantly higher quality, sharper textures, and more accurate geometric details.
Text-to-3D Generation We apply the multi-view diffusion models from Instant3D (Li et al., 2024) to generate 4-view images from text inputs, followed by our Mesh-LRM to achieve text-to-3D generation. Our results and comparison with the original Instant3D pipeline (In3D-LRM) (quantitatively compared in Tab. 5 in terms of reconstruction) are shown in the Appendix Fig. 7. As can be seen, our approach leads to better mesh quality and fewer rendering artifacts.
6 Conclusion
We present MeshLRM, a novel Large Reconstruction Model that directly outputs high-quality meshes. Our model leverages the Differentiable Marching Cubes (DiffMC) method and differentiable rasterization to fine-tune a pre-trained NeRF-based LRM, trained with volume rendering. As DiffMC requires backbone efficiency, we propose multiple improvements (tiny shared MLP and simplified image tokenization) to the LRM architecture, facilitating the training of both NeRF and mesh targeting models. We also find that a low-to-high-res training strategy significantly accelerates the training of the NeRF-based model. Compared with previous work, our method provides both quality and speed improvements and generates high-quality meshes. Finally, we show that our method can be directly applied to applications such as text-to-3D and image-to-3D generation. As meshes are the most widespread format for 3D assets in the industry, we believe our method takes a step towards better and faster integration of 3D asset creation networks in existing 3D workflows and manipulation tools.
Appendix
Appendix A Different between ray opacity loss and other regularizations
The main difference between our ray opacity loss and previous approaches lies in the selection of supervision points. Standard density regularization uniformly supervises all points in space, including those on the mesh surface as well as in empty regions, which can lead to gradients that are less effective at guiding the model toward a clean surface. On the other hand, standard depth loss only supervises points on the mesh surface; once floaters appear during training, its gradients can move these floaters but cannot easily eliminate them. In contrast to these standard losses, our opacity loss supervises the entire space between the surface and the camera, effectively preventing floaters and stabilizing the training process.
Appendix B Comparison with per-scene optimization methods